OpenStudy (anonymous):

Vicky records the low temperature for three consecutive days. Between the first and second days the temperature increased 6 degrees. Between the second and third days the temperature decreased 3 degrees. On the third day the low temperature was –6º. What was the low temperature on the first day?

OpenStudy (gamer56):

i will not feed you the answer this is open study lol so first tell me how you would do this problem?

OpenStudy (gamer56):

if you do not respond to me i can not help you

OpenStudy (anonymous):

I didn't notice anyone typing, I was a different website. Plus, I don't know what to do. Thats why I put it up so someone can help/teach me.

OpenStudy (gamer56):

oh ok here is a link that will answer your questions and give you some answers http://answers.yahoo.com/question/index?qid=20130428135655AA71Utj please vote me best answer since i helped and researched for you

OpenStudy (anonymous):

Eh, you seemed pretty cool anyway... plus you were the only one who TRIED to answer my question very quickly. Thank you! :)

OpenStudy (anonymous):

x + 6 - 3 = -6 (solve for x...the first day low temperature)

OpenStudy (anonymous):

But its a multiple choice question, –15 degrees –9 degrees –3 degrees 3 degrees These are all of my choices. Sorry, for not making that clear. :(

OpenStudy (anonymous):

work out the equation and solve for x ... x + 6 - 3 = -6 can you do that ?

OpenStudy (anonymous):

Nope, I got to choose from –15 degrees –9 degrees –3 degrees 3 degrees

OpenStudy (anonymous):

you will find your answer by working out the equation

OpenStudy (anonymous):

I tried working out the equation, but I couldn't figure it out. That's why I posted the question for help. But, it's apparent we both don't know. Thank you for your time and trying to help! :) Ps. I'm going to post another question... put anything down('i don't care what it is) I'll give you a badge! :)

OpenStudy (anonymous):

x - 6 + 3 = -6 x + 3 = -6 x = -6 - 3 x = -9 degrees

OpenStudy (anonymous):

THANKS! Hope you're right! I'm going to trust you!:)

OpenStudy (anonymous):

wait.... x + 6 - 3 = -6 x + 3 = -6 x = -6 - 3 x = -9 I wrote it down wrong earlier, but this is the correct way....answer is -9 degrees

OpenStudy (anonymous):

Okey, dokey!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

;)